OpenStudy (gabylovesyou):

A theme park company is opening a marine-inspired park in your city. They are in the process of designing the theatre where a killer whale show will take place. The following design is already under construction and will house the whales that perform in the show:

OpenStudy (gabylovesyou):

Main Show Tank Calculation: The main tank has a radius of 70 feet. What is the volume of the quarter-sphered sized tank? Round your answer to the nearest whole number. You must explain your answer using words, and you must show all work and calculations to receive credit.

OpenStudy (gabylovesyou):

@jim_thompson5910 @ganeshie8 @dumbcow

OpenStudy (mathmath333):

\(\large\tt \color{black}{volume~~of~~full~~sphere(v)=\dfrac{4\pi r^3}{3}}\) \(\large\tt \color{black}{as~~the~~quarter~-~sphere~~is~~asked~~(v')=\dfrac{v}{4}}\) \(\large\tt \color{black}{(v')=\dfrac{4\pi r^3}{4\times3}}\)

OpenStudy (mathmath333):

\(\large\tt \color{black}{(v')=\dfrac{\pi r^3}{3}}\)

OpenStudy (gabylovesyou):

@jim_thompson5910 im confused

OpenStudy (anonymous):

Use the equation that mathmath333 gave you. You just need to plug in the radius and solve for v'

OpenStudy (anonymous):

The equation would look like \[V'=\frac{ \pi(70)^{3} }{ 3 }\]